# Affine Chord Sobolev Inequalities from Integral Geometry

### v1

@@ 9.2. Results for $s$-concave functions
Let $g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$ with fixed $u\in\mathbb{S}^{n-1}$. By Theorem 9.8, letting $\omega(t)=g(0)(1-s^{\prime}t)_{+}^{1/s^{\prime}}$ and $\varphi(t)=\frac{1}{s^{\prime}}(1-(g(tu)/g(0))^{s^{\prime}})$ in Lemma 9.2, we obtain
@@
where $s^{\prime}=s/(ns+1)$. Then $c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f$ implies that $\zeta(\alpha)=\zeta(\beta)$. Since $\omega$ is strictly decreasing on its support, by Lemma 9.3, we obtain $\lambda=\zeta(\alpha)^{-1}$ such that
$\varphi(t)=\frac{1}{s^{\prime}}\Big(1-\Big(\frac{g(tu)}{g(0)}\Big)^{s^{\prime}}\Big)=\lambda t.$
This implies that $g(tu)^{s^{\prime}}$ is linear in $t$. Then we conclude the proof by Lemma 9.10. ∎

### v2

@@ 6.2. Results for $s$-concave functions
Let $g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$ for a fixed $u\in\mathbb{S}^{n-1}$ and $r\in\mathbb{R}$. Setting $\omega(r)=g(o)(1-s^{\prime}r)_{+}^{1/s^{\prime}}$ and $\varphi(r)=\frac{1}{s^{\prime}}(1-(g(ru)/g(o))^{s^{\prime}})$ in Lemma 6.2, we obtain
@@
where $s^{\prime}=s/(ns+1)$. Then, the equality $c_{n,\beta}(s){\rm R}_{\beta}f=c_{n,\alpha}(s){\rm R}_{\alpha}f$ implies that $\zeta(\alpha)=\zeta(\beta)$. Since $\omega$ is strictly decreasing on its support, Lemma 6.3 yields $\lambda=\zeta(\alpha)^{-1}$ such that
$\varphi(r)=\frac{1}{s^{\prime}}\Big(1-\Big(\frac{g(ru)}{g(o)}\Big)^{s^{\prime}}\Big)=\lambda t.$
This shows that $g(ru)^{s^{\prime}}$ is affine in $r$, and the conclusion follows from Lemma 6.15. ∎
We conclude this section by showing the Rogers-Shephard inequality and Zhang inequality for $s$-concave functions, which are special cases of Theorem 3.
Theorem 6.17 (Rogers-Shephard inequality for $s$-concave functions).
Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. Then
$|{\rm D}\,{\rm supp}(f)|\leq\frac{\Pi_{k=1}^{n}(k+\frac{1}{s})}{n\cdot n!\|f\|_{1}}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy.$
Equality holds if and only if $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $x_{0}$ is the maximizer of $f$ with $a=f(x_{0})=\|f\|_{\infty}$, and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing $x_{0}$.
Theorem 6.18 (Zhang’s inequality for $s$-concave functions).
Let $s>0$ and $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, non-negative, $s$-concave function with compact support. Then
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\min\{f(x),f(y)\}dxdy\leq\frac{n\cdot n!(n+\frac{1}{s})^{n}}{\Pi_{k=1}^{n}(k+\frac{1}{s})}\|f\|_{1}^{n+1}|\Pi^{*}f|.$
Equality holds if and only if $f(x)=a(1-\|x-x_{0}\|_{\triangle-x_{0}})_{+}^{1/s}$, where $x_{0}$ is the maximizer of $f$ with $a=f(x_{0})=\|f\|_{\infty}$, and $\triangle$ is a simplex in $\mathbb{R}^{n}$ containing $x_{0}$.